OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Write your answer as a mixed number. Base 3 1/2 feet, altitude 3/4 feet. Area = _____sq. ft.

Directrix (directrix):

A = bh A = (3 1/2) * (3/4) A = (7/2) * (3/4) A = (7*3)/(2*4) A = 14/12 = 7/6 = 1 1/6 square feet

OpenStudy (anonymous):

when i type my answer you gave me in the computer it says its wrong. i dont understand

OpenStudy (anonymous):

it should be 21/8 = 2 5/8.

OpenStudy (anonymous):

yes thats right thx!!

Directrix (directrix):

Sorry for my error in arithmetic A = bh A = (3 1/2) * (3/4) A = (7/2) * (3/4) A = (21/8) A = 2 2/5 square feet